# From Tokens to Thoughts: How LLMs and Humans Trade Compression for Meaning

### v4

@@ 4.2 The $\mathcal{L}$ Objective: Balancing Representational Complexity and Semantic Distortion
$\mathcal{L}(X,C;\beta)=\text{Complexity}(X,C)+\beta\cdot\text{Distortion}(X,C).$ (1)
Here, $\beta\geq 0$ is a hyperparameter that balances the relative importance of the two terms.
The Complexity (Rate) Term: The first component, Complexity($X,C$), measures the informational cost or intricacy of representing the original items $X$ through their assignments to clusters $C$. It is quantified by the mutual information $I(X;C)$ between the items and their cluster labels. A lower $I(X;C)$ signifies greater compression, meaning the cluster assignments $C$ make the specific items $X$ more predictable (i.e., require less information to specify beyond the cluster label). Defining I(X;C)=H(X)H(X|C)I(X;C)=H(X)-H(X|C), and assuming $|X|$ equiprobable unique items for the initial entropy calculation ($H(X)=\log_{2}|X|$), the conditional entropy is $H(X|C)=\frac{1}{|X|}\sum_{c\in C}|C_{c}|\log_{2}|C_{c}|$. This assumes that for this complexity calculation, items within each cluster $C_{c}$ (of size $|C_{c}|$) are indistinguishable beyond their shared label $c$. Thus:
$\text{Complexity}(X,C)=\log_{2}|X|-\frac{1}{|X|}\sum_{c\in C}|C_{c}|\log_{2}|C_{c}|.$ (2)
This term formalizes the representational compactness aspect central to [RQ1].
The Distortion Term: The second component, Distortion($X,C$), quantifies the loss of semantic fidelity incurred by grouping items into clusters. It is measured as the average intra-cluster variance of the item embeddings, reflecting how tightly items are bound to their cluster’s central tendency and thus the cluster’s semantic coherence. This directly relates to the preservation of fine-grained semantic information, an idea explored in [RQ2]. For each cluster $c\in C$, its centroid is $x_{c}=\frac{1}{|C_{c}|}\sum_{x\in c}x$ (the mean embedding of its items). Its internal variance is $\sigma_{c}^{2}=\frac{1}{|C_{c}|}\sum_{x\in c}\|x-x_{c}\|^{2}$. The total distortion for the clustering $C$ is the weighted average of these variances:
$\text{Distortion}(X,C)=\frac{1}{|X|}\sum_{c\in C}|C_{c}|\cdot\sigma_{c}^{2}.$ (3)
A lower distortion value implies that, on average, items are close to their respective cluster centroids, suggesting better preservation of shared semantic features within each cluster.
The Unified Objective Function: Substituting the formal definitions of Complexity (Equation 2) and Distortion (Equation 3) into our general formulation for $\mathcal{L}$ (Equation 1) yields the complete objective function that underpins our comparative analysis:
$\mathcal{L}(X,C;\beta)=\left(\log_{2}|X|-\frac{1}{|X|}\sum_{c\in C}|C_{c}|\log_{2}|C_{c}|\right)+\beta\cdot\left(\frac{1}{|X|}\sum_{c\in C}|C_{c}|\cdot\sigma_{c}^{2}\right).$ (4)
This $\mathcal{L}$ function provides a single, principled measure for evaluating how effectively a given clustering $C$ balances the need for informational compression against the imperative to preserve semantic meaning, serving as the direct quantitative tool for addressing [RQ3].
With the $\mathcal{L}$ objective now fully specified, our information-theoretic framework provides a comprehensive toolkit. The Complexity term (Equation 2) allows us to quantify aspects of representational compactness pertinent to [RQ1], while the Distortion term (Equation 3) enables the assessment of semantic preservation, crucial for [RQ2]. The overall $\mathcal{L}$ function (Equation 4) then directly facilitates the evaluation of the integrated compression-meaning trade-off, central to [RQ3]. Thus, this framework equips us to systematically and quantitatively investigate how LLMs and human cognition manage the balance between informational efficiency and semantic richness. We apply this framework in our empirical investigation detailed in Section 5.
5 Unpacking Representational Strategies: An Empirical Investigation
Building on our information-theoretic framework (Section 4) and established benchmarks (Section 3), we now empirically investigate our research questions. This section details the specific methodologies employed to compare LLM and human conceptual strategies across the key dimensions of conceptual alignment, internal semantic structure, and overall representational efficiency.
[RQ1] Assessing Conceptual Alignment To investigate how LLM-derived conceptual categories align with human-defined ones (RQ1), probing representational compactness, we cluster LLM token embeddings using k-means ($K$ set by human category counts per dataset). Alignment with human categories is quantified using Adjusted Mutual Information (AMI), Normalized Mutual Information (NMI), and Adjusted Rand Index (ARI), against a random clustering baseline.
[RQ2] Examining Internal Cluster Geometry and Semantic Preservation To assess how LLM representations capture human-like typicality (RQ2), examining internal category geometry, we calculate the cosine similarity of each item’s token embedding to the token embedding of its human-assigned category name (e.g., ‘robin‘ to ‘bird‘). These LLM-derived similarities are then correlated (Spearman’s $\rho$) with human typicality ratings from our cognitive science datasets.
[RQ3] Evaluating the Efficiency of the Compression-Meaning Trade-off To evaluate the overall balance of compression and meaning (RQ3), we apply our framework by computing the $\mathcal{L}$ objective (Equation 4, $\beta=1$) for both human and LLM-derived conceptual structures (the latter from k-means over a range of $K$). This compares how each system balances Complexity $I(X;C)$ against Distortion. Cluster entropy is an ancillary measure of compactness.
For robustness, all k-means clustering involves one hundred random initializations with averaged results. Appendix A.8 provides details on supplementary metrics like Silhouette scores.
@@ 5 Unpacking Representational Strategies: An Empirical Investigation
We first investigate whether LLMs form conceptual categories aligned with human judgment.
@@
Key Finding: Broad Alignment with Human Categories LLM-derived clusters significantly align with human-defined conceptual categories, suggesting they capture key aspects of human conceptual organization. Notably, certain encoder models exhibit surprisingly strong alignment, sometimes outperforming much larger models, highlighting that factors beyond sheer scale influence human-like categorical abstraction.
Experimental Recap: LLM token embeddings from our benchmark datasets (Rosch, 1973a, 1975; McCloskey and Glucksberg, 1978) were clustered (k-means; $K$ matching human category counts). Alignment with human categories was measured using AMI, NMI, and ARI (AMI shown in Figure 1; see Appendices A.8, A.9 for full details). Both the static and the contextual embeddings were extracted. The contextual embeddings were measured layer-wise, and reported using Peak AMI across the layers. Note that GloVe and Word2Vec only have static embeddings.
Results and Observations: Across all tested LLMs, derived conceptual clusters aligned with human categories significantly above random chance (Figure 1, showing averaged peak AMI scores). This indicates their semantic spaces encode information supporting human-like grouping at a macro level. Notably, the BERT family (especially BERT-large-uncased) demonstrated robust alignment, often comparable to or exceeding that of much larger decoder-only models. This suggests that architectural or pre-training factors, not just scale, influence the formation of human-like categorical structures.
Figure 1: LLM-derived Clusters Show Above-Chance Alignment with Human Conceptual Categories. Adjusted Mutual Information (AMI) between human categories and LLM-embedding clusters versus model size. Results are averaged over three psychological datasets. All models perform significantly better than random clustering. Encoder architectures achieve consistently high performance across modalities and scales; notably, the classic Word2Vec and GloVe models, despite offering only static, outdated embeddings, reach static AMI scores that match the peak AMI of modern LLMs. See Appendix A.4 for a comparison of static, mean, and peak AMI.
Interpretation: These findings confirm that LLMs can recover broad, human-like categories from their embeddings, validating deeper comparative analyses. This macro-level agreement necessitates examining the finer-grained internal geometry of these categories, which we address next.
Influence of Training Dynamics on Alignment: To explore how semantic structure develops during training, we applied both the static embedding and the contextual (layer-wise) analysis on 10% of OLMo’s checkpoints (spaced evenly across training steps). We find that concept representations improve steadily during training, with a rapid gain in structure early on, followed by a slower, more gradual refinement. Interestingly, the model exhibits a two-phase dynamic: initial optimization moves semantic processing from deep layers to mid-network layers during rapid learning, followed by continued architectural reorganization that optimizes efficiency while maintaining performance (Appendix A.5 Figure 8). This suggests that concepts emerge early in training but are progressively consolidated and reallocated across layers, reflecting a shift from brute-force memorization toward more efficient, distributed representations. Such dynamics highlight that alignment is not just a final-state property of the trained model, but is deeply shaped by the trajectory of learning itself.
5.2 [RQ2] Delving Deeper: Fidelity to Fine-Grained Semantics
Having established that LLMs broadly align with human conceptual categories (Section 5.1), we next investigate a more nuanced question: Do LLMs also capture the internal semantic structure of these categories, particularly human-like item typicality?
@@
Experimental Recap: For this RQ, as detailed in this section’s introduction, we compared human typicality judgments from the cognitive science datasets (Rosch, 1973a, 1975; McCloskey and Glucksberg, 1978) with an LLM-based measure. Specifically, we calculated the cosine similarity between each item’s token embedding and the token embedding of its human-assigned category name (e.g., ‘robin‘ vs. ‘bird‘). These item-to-category-label similarities were then correlated (Spearman’s $\rho$ (Wissler, 1905)) with human-rated typicality scores.
@@
We employed two analysis approaches: static-layer analysis utilized static items/category embeddings from the embedding layer, while peak AMI layer analysis utilized items/category contextual embeddings from the layer that maximized AMI as measured in RQ1. To extract category embeddings in the peak AMI approach, we replaced the item with the category name in the same prompt template (see template and template- and pooling robustness analysis in Appendix A.4).
Results and Observations: Spearman correlations between LLM-derived item-to-category-label similarities and human typicality judgments are generally modest across most models and datasets, with peak AMI layers showing particularly weak correlations (Tables 3,  4 in Appendix A.10; Figure 15). Even when correlations reach statistical significance ($p<0.05$), their magnitudes typically indicate a limited correspondence.
@@
Several architectural patterns emerge: representation-focused models (Word2Vec, GloVe) and some encoder models (T5-large, the two ViT encoders, and BERT large) demonstrate stronger static-layer performance than decoder-only models (Llama, Gemma, etc.), suggesting that models trained specifically for representation learning may be more effective for capturing semantic category relationships in their embeddings. In contrast, modern decoder-only models show consistently lower static-layer correlations, potentially due to their training objectives being more focused on next-token prediction rather than semantic representation.
Notably, peak AMI layer analysis shows that correlations weaken substantially relative to static-layer analysis across most architectural variants, suggesting that cross-cluster relations (RQ1) and within-cluster organization (RQ2) are encoded at different layers. This implies that downstream applications should select layers based on the resolution they require; whether capturing broad cross-cluster structure or fine-grained within-cluster organization. Furthermore, while some architectural approaches capture certain aspects of human typicality judgments, their internal representations exhibit only modest alignment with human-perceived fine-grained semantic distinctions, such as item typicality or psychological distance to category prototypes. As a result, no single architectural approach robustly reproduces the full spectrum of human typicality gradients using this measure. Additional visualizations supporting these findings are provided in Appendix A.11.
Interpretation: These findings suggest that while LLMs can identify features for broad categorization, their organization of semantic space around explicit category labels does not fully mirror the nuanced prototype structures evident in human typicality judgments. The factors driving an item’s embedding similarity to its category label’s embedding in LLMs may differ from the rich, multifaceted criteria (e.g., perceptual attributes, functional roles) underpinning human typicality. LLMs might instead capture a more statistically uniform association to category labels, thereby under-representing the graded, prototype-centric nature of human concepts. The tension between representational compactness and semantic fidelity becomes particularly evident when comparing peak AMI layers to static layers: while peak AMI layers optimize the Information Bottleneck trade-off for efficient clustering, they may prioritize compression over preserving the nuanced semantic information needed for human-like typicality judgments. This divergence in capturing fine-grained semantics leads to our subsequent inquiry into overall information processing efficiency.
@@
Having explored categorical alignment (RQ1) and internal semantic structure (RQ2), we now address our central question: How do LLM and human representational strategies compare in overall efficiency when balancing informational compression against semantic meaning preservation? Our information-theoretic framework directly probes this trade-off.
@@
Key Finding: Divergent Efficiency Strategies LLMs demonstrate markedly superior information-theoretic efficiency in their conceptual representations compared to human conceptual structures. Evaluated via our $\mathcal{L}$ objective, LLM-derived clusters consistently achieve a more “optimal“ balance (by this measure) between representational complexity (compression) and semantic distortion. Human conceptualizations, while richer, appear less statistically compact, suggesting optimization for pressures beyond pure statistical compressibility.
@@
Experimental Recap: As detailed in this section’s introduction, we analyzed human-defined categories and LLM-derived clusters (from k-means across various $K$) using two primary information-theoretic measures: mean cluster entropy ($S_{\alpha}$) (Giraldo et al., 2014; Wei et al., 2025) and our $\mathcal{L}$ objective function (Equation 4, with $\beta=1$).
Results and Observations: Illustrative results from one dataset (Rosch, 1975) are shown in Figure 2; trends were consistent across all datasets (full results in Appendix A.13).
@@
Cluster Entropy Insights: Human concepts consistently exhibit higher mean entropy than LLM-derived clusters, even at similar $K$ values (Figure 2, left). This suggests that, by this measure, human categories are less statistically “compact“ and encompass greater internal diversity than LLM clusters.
(a) Distortion as a function of Complexity demonstrates the balance of the Information Bottleneck in our information-theoretic objective ($\mathcal{L}$). The plot illustrates how the IB balances compression (complexity) against information preservation (distortion). Encoder models (BERT, ViT, classic models) show lower distortion rates compared to all the decoder models across all complexity values.
(b) Contextual embeddings achieve better-than-human $\mathcal{L}$ trade-off. Our information-theoretic objective ($\mathcal{L}$) vs. $K$. Lower $\mathcal{L}$ indicates a more statistically optimal compression-meaning balance compared to human baseline.
Figure 2: LLMs Show More Statistically “Optimal” Compression Than Humans in the $\mathcal{L}$ Measure. (a) Distortion as a function of Complexity showing the trade-off relationship between the two terms in the IB-RDT objective ($\mathcal{L}$). Encoder models (BERT, ViT, classic models) show lower distortion rates compared to all the decoder models across all complexity values. (b) IB-RDT objective ($\mathcal{L}$) as a function of $K$ used for k-means clustering using contextual embeddings. Lower $\mathcal{L}$ indicates a more statistically optimal compression-meaning balance. Human categories consistently show higher $\mathcal{L}$ values. Results shown for Rosch (1975) dataset; full results in Appendix A.13.
@@
Information-Theoretic Objective ($\mathcal{L}$) Insights: The $\mathcal{L}$ objective reveals an even starker divergence (Figure 2, right). LLM-derived clusters consistently achieve significantly lower $\mathcal{L}$ values than human conceptual categories across most tested $K$. Since a lower $\mathcal{L}$ signifies a more statistically “optimal” trade-off between minimizing complexity and distortion within our framework, this implies LLMs are more “efficient” by this specific information-theoretic benchmark.
Interpretation: The combined results from entropy and the $\mathcal{L}$ objective strongly indicate a fundamental difference in representational strategy. LLMs appear highly optimized for statistical compactness, achieving information-theoretically “efficient” representations by minimizing redundancy and internal variance. Human conceptual systems, in contrast, while appearing “suboptimal“ by these statistical measures, are likely shaped by a broader array of functional imperatives. These include the demands of adaptive generalization, rich causal and functional inference, the constraints of neural embodiment, and the requirements of nuanced communication–pressures that may favor representations less statistically “tidy” but ultimately more flexible and powerful for navigating a complex world.
The Objective $\mathcal{L}$ vs. Downstream Task Performance: The relationship between concept compression, as measured by the $\mathcal{L}$ objective function, and downstream task performance shows no correlation, indicating that concept compression captures an orthogonal property to benchmark accuracy. See Appendix A.14 for detailed analysis.

### v5

@@ 4.2 The $\mathcal{L}$ Objective: Quantifying the Trade-off
$\mathcal{L}(X,C;\beta)=\underbrace{I(X;C)}_{\text{Complexity: bits needed}}+\beta\cdot\underbrace{\frac{1}{|X|}\sum_{c\in C}\sum_{e_{i}\in c}\|e_{i}-\bar{e}_{c}\|^{2}}_{\text{Distortion: semantic spread}}$ (3)
where $\beta$ weights the relative importance of compression versus coherence.
4.2.1 The Complexity Term: Measuring Compression
Complexity quantifies how much information the clustering preserves about individual items through mutual information $I(X;C)$. Intuitively, if knowing an item’s cluster tells us little about which specific item it is, compression is high (low complexity).
Given $|X|$ items partitioned into clusters of sizes $\{|C_{c}|\}$:
$\text{Complexity}(X,C)=I(X;C)=\log_{2}|X|-\frac{1}{|X|}\sum_{c\in C}|C_{c}|\log_{2}|C_{c}|$ (4)
This equals the reduction in uncertainty about item identity when told its cluster. Uniform clusters minimize complexity (one $|X|$ cluster; maximum compression), while singleton clusters maximize it (no compression).
4.2.2 The Distortion Term: Measuring Semantic Coherence
Distortion captures how well clusters preserve semantic relationships and meanings by measuring the average squared distance between items and their cluster centroids in embedding space (spread):
$\text{Distortion}(X,C)=\frac{1}{|X|}\sum_{c\in C}|C_{c}|\cdot\sigma^{2}_{c}$ (5)
where $\sigma^{2}_{c}=\frac{1}{|C_{c}|}\sum_{e_{i}\in c}\|e_{i}-\bar{e}_{c}\|^{2}$ is the variance within cluster $c$, and $\bar{e}_{c}$ is its centroid.
Low distortion indicates tight, semantically coherent clusters with similar embeddings. This geometric measure directly captures what we intuitively mean by “meaningful” categories: robins and sparrows cluster tightly as similar birds, while bats would increase distortion on downstream tasks if categorized with them.
4.3 Connecting Framework to Research Questions
Our framework provides unified metrics for all three research questions:
[RQ1] Categorical Alignment: How do LLMs and humans partition semantic space? The Complexity term $I(X;C)$ directly measures this by quantifies how many bits are needed to specify individual items given their clusters. Comparing $I(X;C_{\text{Human}})$ with $I(X;C_{\text{LLM}})$ reveals whether both systems create similarly-sized groupings with comparable compression rates. Higher mutual information means finer-grained categories; lower means broader, more compressed groupings.
[RQ2] Internal Semantic Structure: Do LLMs capture human-like typicality gradients? The Distortion term measures how well clusters preserve semantic coherence. In our typicality analysis, we check whether typical items cluster tightly near centroids while atypical items lie farther away. Low distortion with clear center-periphery structure indicates prototype organization that mirrors human cognitive structure.
[RQ3] Compression-Meaning Trade-off: How do different systems balance efficiency against semantic fidelity? The complete $\mathcal{L}$ objective reveals fundamental optimization strategies. By varying $K$ (number of clusters) and computing $\mathcal{L}$ curves, we uncover system priorities: aggressive compressors rapidly achieve low $\mathcal{L}$ values by sacrificing nuance, while systems preserving semantic richness maintain higher $\mathcal{L}$ to retain meaningful distinctions. The shape and level of these curves expose whether a system optimizes for statistical efficiency or cognitive utility.
5 An Empirical Investigation of Representational Strategies
Building on our information-theoretic framework (Section 4) and established benchmarks (Section 3.1), we empirically investigate how LLMs and humans navigate the compression-meaning trade-off. For each analysis, we examine both static embeddings and contextual embeddings (across all hidden layers), revealing how and when context shapes conceptual organization.
@@ 5 An Empirical Investigation of Representational Strategies
We first investigate whether LLMs form conceptual categories aligned with humans, which examines how information is compressed into discrete groups (the complexity term in our framework).
@@
Key Finding: Broad Alignment with Human Categories LLM-derived clusters significantly align with human-defined conceptual categories, suggesting they capture key aspects of human conceptual organization. Surprisingly certain encoder models exhibit strong alignment, sometimes outperforming much larger models, highlighting that factors beyond sheer scale influence human-like categorical abstraction.
Approach: We tested whether LLMs naturally organize our 1,049 items into categories resembling human conceptual structure. Token embeddings were extracted at two levels: (i) static embeddings from input layers (E matrix), representing context-free lexical knowledge; (ii) contextual embeddings from all hidden layers, measured layer-wise to identify peak conceptual alignment. These were clustered using k-means ($K$ matching human category counts) and evaluated against human categories using Adjusted Mutual Information (AMI), Normalized Mutual Information (NMI), and Adjusted Rand Index (ARI) metrics.
Broad Categorical Agreement: All 40+ models achieve significant above-chance alignment (Figure 1 (Left)). Even at baseline, static embeddings show substantial alignment (mean $AMI\approx 0.45$), which contextual processing enhances to peak $AMI\approx 0.55$. This confirms that LLMs encode human-like categorical boundaries. Full NMI and ARI results in Appendices B.8B.9.
Architecture Matters More Than Scale: Surprisingly, BERT-large-uncased (340M parameters) achieves $AMI=0.60$, matching or exceeding models 100× larger. Classic Word2Vec and GloVe, despite having only static embeddings and predating modern architectures by years, reach AMI scores rivaling contemporary LLMs’ peak performance. This suggests that fundamental semantic structure emerges from relatively simple distributional learning, with encoder architectures particularly effective at capturing human-like categories regardless of scale.
Figure 1: LLMs capture categorical boundaries but miss semantic gradients. Left: All 40+ models achieve above-chance AMI with human categories, with encoder architectures (squares) matching or exceeding decoder models 100× larger. Results show the layer with peak AMI score per model, see static and mean scores in Figure A.7. Right: Despite categorical success, models show weak correlations ($\rho<0.2$ for most) with human typicality judgments, revealing divergent representational strategies. This divergence between capturing boundaries (compression) while missing internal structure (meaning) reveals how LLMs and humans fundamentally differ in their representational strategies. We note that encoder models align more than decoder models, but these correlations are still modest. Computed using the static embeddings, see full results in Tables 3-4 in Appendix B.10.
5.2 [RQ2] Zooming In: Fidelity to Fine-Grained Semantics
Having established broad categorical alignment, we now examine whether LLMs capture the internal semantic structure of categories. Specifically, we check how meaning is preserved within clusters.
@@
Approach: We test whether LLMs encode human-like typicality gradients, i.e., whether robins are more “birdy” than penguins. For each item, we compute the cosine similarity with its category name using embeddings (e.g., ‘robin’ → ‘bird’). We then compare these similarities with human typicality ratings using Spearman’s correlation coefficient $\rho$ (Wissler, 1905).
@@
We employed two analysis approaches: (i) static-layer analysis using embeddings directly from the input layer; (ii) peak AMI layer analysis using contextual embeddings from the layer that maximized AMI in RQ1. For the peak AMI approach, we extracted category embeddings by replacing items with category names in the same prompt template, ensuring consistent contextualization (see Appendix B.4 for template and pooling robustness analysis).
Weak Typicality Alignment: Correlations between LLM internal organization of concepts and human typicality are modest at best (Figure 1 (Right); Tables 3-4 in Appendix B.10). Static embeddings show weak correlations: BERT achieves $\rho=0.38$ ($p<0.05$), while most decoder models fall below $\rho=0.15$. Even when statistically significant, these correlations indicate limited correspondence with human judgments. This shows that the internal concept geometries of models differ from those of humans, with representation-focused models aligning more closely.
Architectural Patterns: Several clear patterns emerge in how different architectures capture typicality. Representation-focused models (Word2Vec, GloVe) and most encoder models (both ViT encoders, BERT-large) demonstrate stronger static-layer performance than decoder-only models (Llama, Gemma, Qwen families). Static correlations range from $\rho\approx 0.25$-$0.40$ for representation-focused models versus $\rho<0.15$ for most decoders.
@@
This divergence likely stems from training objectives: models explicitly trained for representation learning appear more effective at capturing semantic category relationships in their embeddings, while modern decoder-only models, which optimized primarily for next-token prediction, show consistently lower static-layer correlations. The pattern holds across model scales, suggesting architectural design matters more than size for capturing semantic gradients.
Layer-wise Analysis Reveals a Trade-off: Comparing static and peak AMI layers exposes an architectural limitation. Peak AMI layers, which are optimal for clustering, show systematically weaker typicality correlations than static layers. This pattern holds across model families: layers that best separate categories (RQ1) poorly preserve within-category structure (RQ2). The implication is clear: current architectures encode different aspects of meaning at different depths, forcing applications to choose between broad categorization and semantic nuance.
Interpretation: The divergence between LLMs and humans reflects fundamentally different organizational principles. Humans judge typicality through rich, multidimensional criteria: robins are typical birds due to size, flight ability, song, frequency of encounter, etc. This creates graded categories with clear prototypes and cognitive structures that optimize flexible reasoning and generalization.
LLMs, in contrast, appear to encode flatter statistical associations between items and category labels. Although sufficient for categorization and fluent text generation, these representations miss the prototype structure that makes categories cognitively useful. This difference suggests that LLMs optimize for different objectives than human cognition, a hypothesis that we test directly in RQ3 by examining how each system balances compression against semantic preservation.
@@
Having explored categorical alignment (RQ1) and internal semantic structure (RQ2), we now address our central question: How do LLM and human representational strategies compare when balancing compression against meaning preservation?
@@
Key Finding: Divergent Efficiency Strategies LLMs demonstrate markedly superior information-theoretic efficiency compared to human conceptual structures. Evaluated via our $\mathcal{L}$ objective, LLM-derived clusters consistently achieve more “optimal” compression-meaning balance. Human conceptualizations, while richer, appear less statistically compact, suggesting optimization for cognitive flexibility over pure statistical efficiency.
@@
Approach: We analyzed human-defined categories and LLM-derived clusters using our $\mathcal{L}$ objective function (Equation 3, $\beta=1$) and mean cluster entropy ($S_{\alpha}$). For LLMs, we performed k-means clustering across various $K$ values to trace the full compression-meaning frontier.
Results: Our analysis reveals three key patterns (Figure 2; full results in Appendix B.13):
LLMs Achieve Superior Statistical Efficiency: Both measures reveal stark differences between LLMs and humans (Figure 2; full results in Appendix B.13):
@@
Higher human entropy. Human concepts consistently exhibit higher mean cluster entropy than LLM-derived clusters at comparable $K$ values, indicating less statistical compactness but greater internal diversity.
@@
Lower LLM $\mathcal{L}$ scores. LLM-derived clusters achieve significantly lower $\mathcal{L}$ values than human categories across all tested $K$ (Figure 2b). Since lower $\mathcal{L}$ signifies more optimal compression-distortion balance, LLMs are demonstrably more “efficient” by this information-theoretic measure.
Architectural differences. The Complexity-Distortion plot (Figure 2a) reveals that encoder models (BERT, ViT, classic models) consistently achieve superior trade-offs. Their distortion at any given complexity is lower compared to decoder models. This advantage persists across both static and contextual embeddings.
Statistical Optimality Versus Cognitive Utility: This divergence reveals fundamental differences in optimization pressures. LLMs, trained on massive text corpora, develop maximally efficient statistical representations that minimize redundancy and internal variance. Human conceptual systems, though appearing “suboptimal” by information-theoretic measures, serve distinct cognitive needs: flexible generalization, causal reasoning, and robust communication under neural constraints.
Our analysis reveals that compression efficiency does not predict functional capability. We find no correlation between $\mathcal{L}$ scores and downstream performance (r=0.20r=-0.20, $\rho=0.51$ on MMLU; Appendix B.14). This suggests that apparent human “inefficiency” reflects optimization for cognitive flexibility rather than statistical compression. While LLMs excel at compact representation, they may sacrifice the semantic richness essential for human-like understanding.
The consistent architectural patterns we observe raise fundamental questions: Do understanding and generation require distinct computational strategies? The superiority of representation-focused models suggests that current unified architectures may be conflating two fundamentally different cognitive tasks.
(a) Complexity-distortion trade-off across architectures
(b) $\mathcal{L}$ objective comparison with human categories
Figure 2: Divergent optimization strategies: LLMs achieve superior information-theoretic efficiency while humans preserve semantic richness. (a) Encoder models (BERT, ViT, classic embeddings) consistently achieve lower distortion than decoder models at any given complexity level. (b) All LLM-derived clusters achieve lower $\mathcal{L}$ values than human categories (dashed line), indicating more “optimal” compression-distortion balance. Data from Rosch (1975).
5.4 Emergence During Training: How Divergent Strategies Develop
Having established that LLMs and humans employ divergent representational strategies, we investigate how these strategies emerge. Analysis of OLMo-7B across 57 training checkpoints (1K to 557K steps, approximately 4B to 2.5T tokens) reveals how conceptual structure emerges.
Two-Phase Representational Development. Conceptual organization emerges via two phases (Figure 8). First, rapid concept formation (1K-100K steps) establishes basic categorical structure, with AMI rising from near zero to approximately 0.45, achieving 80% of final alignment within just 10% of training. Second, architectural reorganization (100K-500K steps) systematically migrates semantic processing from deeper layers toward mid-network while AMI continues to gradually improve. This migration from layer 29 to layer 23 occurs without sacrificing categorical alignment, suggesting the model discovers more efficient internal representations.
Architectural Reorganization as Optimization. The upward migration of semantic processing hints that the model discovers increasingly efficient encodings. Early training relies on deep, memorization-heavy representations, but as training progresses, the model shifts to distributed mid-network encoding. This reorganization may explain apparent “emergent” capabilities: they arise not from learning fundamentally new information but from more efficient internal organization of existing knowledge.
Implications. These dynamics demonstrate that the compression-oriented strategy observed in fully trained models develops from the earliest stages of training. The rapid initial alignment followed by efficiency-focused reorganization suggests models are inherently biased toward statistical compression rather than semantic richness. Achieving human-like representations may require not just different final objectives but fundamentally different learning dynamics that actively maintain semantic diversity throughout development.